# Improving Acoustic Scene Classification in Low-Resource Conditions

### v1

@@ I Introduction
To tackle these challenges, we propose DS-FlexiNet, which integrates depthwise separable convolutions from MobileNetV2 [8] with ResNet-like residual connections. Quantization Aware Training (QAT) reduces model size and computational requirements without compromising accuracy. Additionally, data augmentation techniques like Auto Device Impulse Response (ADIR) and Freq-MixStyle (FMS) [12, 13] further enhance cross-device generalization. Knowledge distillation from an ensemble of teacher models also improves performance on previously unseen devices.
@@ II-C2 Auto Device Impulse Response
Building on this concept, we propose ADIR Augmentation. By evaluating the energy level of audio samples, we adjust the convolution duration based on the audio’s energy magnitude. For high-energy audio samples, we apply impulse response techniques to preserve intricate characteristics while avoiding distortion. For low-energy samples, we skip impulse response processing to prevent overprocessing and potential distortion. We analyzed energy distribution in the TAU22 dataset (see Fig. 2). ADIR includes a hyperparameter $p_{A}$ to control the probability of DIR convolution.
@@ III Experiment setup
III-A Datset
We use Task 1A data from the TAU Urban Acoustic Scene 2022 Mobile Development dataset [19], comprising 64 hours of recordings from 12 European cities across 10 acoustic scenes and 4 device types: 3 real (A, B, C) and 6 simulated (S1-S6). Additionally, the challenge organizer provides metadata indicating a training set of 13,965 samples and a test set of 2,970 samples.
@@ IV-A Proposed Models
Teacher(w/o) 58.07 70.58 64.90 58.24 56.67 60.38 52.43 58.30 49.36 59.36
Averaged 60.53 71.61 66.50 65.20 59.53 65.78 62.05 61.53 56.71 63.48
Fused 62.33 74.24 68.64 66.44 61.31 67.02 62.71 62.90 59.04 65.26
@@
DS-FlexiNet+FMS+DIR 61.09 71.77 66.43 64.26 60.85 65.72 62.26 62.69 56.88 63.86
DS-FlexiNet+BN+FMS+DIR 58.63 71.90 66.97 64.55 59.91 65.65 60.68 61.63 56.70 63.38
DS-FlexiNet+BN+FMS+ADIR 64.75 72.78 68.54 64.75 62.06 66.23 61.76 62.42 56.45 64.68
DS-FlexiNet+RN+FMS+DIR 63.75 71.78 68.18 65.36 62.48 66.89 61.41 62.92 57.35 64.58
DS-FlexiNet+RN+FMS+ADIR 62.33 74.24 68.64 66.44 61.31 67.02 62.71 62.90 59.04 65.26
@@ IV-C Dimensional Analysis of Residual Normalization
Residual Normalization (RN) significantly improves acoustic scene classification accuracy and robustness across various recording environments and devices. By normalizing along the channel dimension, RN reduces misclassification between similar scenes, such as parks and street traffic, and mitigates channel-specific variations more effectively than frequency normalization. RN achieves up to 86% accuracy in classifying scenes like buses, parks, and street traffic, as shown in Fig. 3(a) and Fig. 3(b).

### v2

@@ I Introduction
To tackle these challenges, we propose DS-FlexiNet, which integrates depthwise separable convolutions from MobileNetV2 [8] with ResNet-like residual connections. Quantization Aware Training (QAT) reduces model size and computational requirements without compromising accuracy. Additionally, data augmentation techniques like Auto Device Impulse Response (ADIR) and Freq-MixStyle (FMS) [12, 13] further enhance cross-device generalization. Knowledge Distillation (KD) from an ensemble of teacher models also improves performance on previously unseen devices.
@@ II-C2 Auto Device Impulse Response
Building on this concept, we propose ADIR method. By evaluating the energy level of audio samples, we adjust the convolution duration based on the audio’s energy level. For high-energy audio samples, we apply impulse response techniques to preserve intricate characteristics while avoiding distortion. For low-energy samples, we skip impulse response processing to prevent overprocessing and potential distortion. We analyzed energy distribution in the TAU22 dataset (see Fig. 2). ADIR includes a hyperparameter $p_{A}$ to control the probability of DIR convolution.
@@ III Experiment setup
III-A Dataset
We use Task 1A data from the TAU Urban Acoustic Scene 2022 Mobile Development dataset [19], comprising 64 hours of recordings from 12 European cities across 10 acoustic scenes and 4 device types: 3 real (A, B, C) and 6 simulated (S1-S6). Additionally, the challenge organizer provides metadata indicating a training set of 139,620 samples and a test set of 29,680 samples.
@@ IV-A Proposed Models
Teacher(w/o) 69.77 64.27 64.90 58.24 56.67 60.38 52.43 58.30 49.36 59.36
Averaged 71.38 62.60 66.50 65.20 59.53 65.78 62.05 61.53 56.71 63.48
Fused 74.00 65.29 68.64 66.44 61.31 67.02 62.71 62.90 59.04 65.26
@@
DS-FlexiNet+FMS+DIR 71.34 64.30 66.43 64.26 60.85 65.72 62.26 62.69 56.88 63.86
DS-FlexiNet+BN+FMS+DIR 70.57 63.76 66.97 64.55 59.91 65.65 60.68 61.63 56.70 63.38
DS-FlexiNet+BN+FMS+ADIR 73.70 66.19 68.54 64.75 62.06 66.23 61.76 62.42 56.45 64.68
DS-FlexiNet+RN+FMS+DIR 72.55 64.12 68.18 65.36 62.48 66.89 61.41 62.92 57.35 64.58
DS-FlexiNet+RN+FMS+ADIR 74.00 65.29 68.64 66.44 61.31 67.02 62.71 62.90 59.04 65.26
@@ IV-C Dimensional Analysis of Residual Normalization
Residual Normalization (RN) significantly improves acoustic scene classification accuracy and robustness across various recording environments and devices. By normalizing along the channel dimension, RN reduces misclassification between similar scenes, such as parks and street traffic, and mitigates channel-specific variations more effectively than frequency normalization. RN achieves up to 86% accuracy in classifying scenes like bus, park, and street traffic, as shown in Fig. 3(a) and Fig. 3(b).